# Guiding an Automatic Speech Recognition Decoder using Large Language Models

### v1

@@ III-C Decoding
The decoding mechanism refers to the process of inferring the most probable sequence of words (or subwords) given an input speech signal. It combines the outputs of the acoustic model and the language model to generate the final transcription using an efficient search algorithm, such as beam search, stack decoding, or fast match [37].
@@
where $\alpha$ scales the influence of the language model, $\beta$ is the word insertion penalty that penalizes (or rewards) the number of words to avoid overgeneration or undergeneration, and $|{W}|$ is the number of words in the hypothesis. This formulation allows a beam search or Viterbi decoder to efficiently explore the space of possible word sequences (as expressed by the HMM states) and choose the one with the highest combined score [33].
In contrast to HMMs, acoustic models that are based on wav2vec 2.0 [9] and HuBERT [10] are trained using the CTC loss function [5] on characters. During decoding of these models, a prefix beam search decoding is used, and the scoring function for each candidate word sequence ${W}$ is formulated as
@@
where we assumed that the posterior probability over the characters $P(Y|X)$ can be easily translated into a probability over words $P_{\text{CTC}}({W}|\mathbf{X})$. Note that systems based on these acoustic models incorporate language models by multiplying the acoustic model $P_{\text{CTC}}({W}|\mathbf{X})$ by $P({W})$ [38, 9]. While this approach yields high performance, it is implausible as it undermines the probabilistic meaning of the models.
@@ IV Proposed Approach
We propose integrating the LLM directly into the decoding process. This iterative approach involves sampling multiple tokens from the LLM at each step and aligning each token with the acoustic model to generate a corresponding acoustic score. As we will explain next, this process employs efficient dynamic programming.
From this point onward, we will use tokens instead of words; thus, $w_{n}\in\mathcal{V}$ denotes a token from the set of all tokens $\mathcal{V}$. Define the alignment sequence $A=(a_{1},\ldots,a_{N})$, where $a_{n}\in[0,T-1]$ represents the start time of token $w_{n}$ for $1\leq n\leq N$. We denote by $\mathcal{A}({W})$ the set of all possible alignments (i.e., start times) corresponding to a given token sequence ${W}$. Lastly, denote by $a_{1}^{n}=(a_{1},\ldots,a_{n})$ the alignment of the first $n$ tokens, $w_{1}^{n}=(w_{1},\ldots,w_{N})$.
Our objective is to derive a MAP estimate of the most likely token sequence, ${W}^{*}$. Rather than incorporating the probability $P({W})$ as a whole, we decompose it into a sequence of token-level predictions. At each step, we evaluate candidate tokens by combining their likelihood under the large language model with an alignment-based acoustic score derived from the emissions of the acoustic model.
More formally, we start from the MAP estimator given in (1):
@@
We approximate this expression by assuming that the most likely alignment dominates the total probability. In other words, we approximate the sum over all alignments by the alignment with the highest probability:
${W}^{*}\simeq\arg\max_{{W}}\max_{A\in\mathcal{A}({W})}P({W},A|\mathbf{X}).$ (9)
This probability can be decomposed into two terms: a term that is associated with the acoustic model and a term that is related to the language model. We start by expressing the sequences ${W}$ and $A$ explicitly:
@@
In summary, we reformulated the MAP estimator as the sum of two terms corresponding to the acoustic and language models. The acoustic model term is expressed as $\sum_{n}P(a_{n}|w_{n},a_{1}^{n-1},w_{1}^{n-1},\mathbf{X})$, in contrast to the conventional forms $P(\mathbf{X}|{W})$ or $P_{\text{CTC}}({W}|\mathbf{X})$ The language model term is given by $\sum_{n}P(w_{n}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$, as opposed to the standard formulation $P({W})=\sum_{n}P(w_{n}|w_{1}^{n-1})$. It is important to note that these terms inherently assume access to the entire speech signal $\mathbf{X}$, which limits the applicability of the proposed method in streaming scenarios. A simple workaround is to process speech in asynchronous chunks, allowing for partial processing. While more sophisticated strategies could be developed to fully enable streaming functionality, such approaches are beyond the scope of this work.
@@ V Algorithmic Implementation
We proceed to show that the MAP estimate in (10) can be evaluated iteratively, allowing the probabilities to be computed incrementally at each step. Suppose we have already computed the probability up to token $n-1$, that is, we know $\log P(w_{1}^{n-1},a_{1}^{n-1}|\mathbf{X})$. The probability of token $n$ is obtained by incrementally adding two terms: the language model probability, $\log P(w_{n}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$ and the acoustic alignment probability $\log P(a_{n}|w_{n},a_{1}^{n-1},w_{1}^{n-1},\mathbf{X})$. Overall, we have,
@@
Practically, the iterative procedure operates as follows: At step $n-1$, we assume that the probability $P(w_{1}^{n-1},a_{1}^{n-1}|\mathbf{X})$ has already been maximized for the optimal token sequence $w_{1}^{n-1}$ and their start-time $a_{1}^{n-1}$ relative to the speech signal. At step $n$, we prompt the LLM with the sequence $w_{1}^{n-1}$ (that were already aligned to the speech) and sample $K$ candidate tokens $\{w_{n}^{(k)}\}_{k=1}^{K}$, each with an associated probability $P(w_{n}^{(k)}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$. Each candidate token is then aligned to the speech signal $\mathbf{X}$, assuming the last token began at $a_{n-1}$. The alignment is performed using the Viterbi algorithm [39], which computes the most likely alignment of each token’s characters to the CTC emissions [^1] For example, this can be implemented by performing forced alignment of the tokens $w_{n}$ to the CTC emission outputs of wav2vec 2.0 or HuBERT.. For each candidate, we evaluate the combined probability of the token and its alignment, namely $P(w_{n}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$ and $P(a_{n}|w_{n},a_{1}^{n-1},w_{1}^{n-1},\mathbf{X})$. The candidate $w_{n}$ and alignment $a_{n}$ with the highest joint probability are selected.
@@
Algorithm 1 Zero-Shot LLM–Driven ASR Decoder
@@
[…]
We conclude this section by summarizing the proposed method in the form of pseudocode, presented in Algorithm 1. The algorithm takes as input a speech signal $\mathbf{X}$, an acoustic model (wav2vec 2.0 [9] or HuBERT [10]) denoted as AM, and a large language model, (e.g., LLaMA 2, Falcon or GPT-2), denoted as LLM. Additional input parameters include the beam size $B$, the number of candidate tokens sampled at each iteration from the LLM $K$, the LLM scaling factor $\alpha$, and the token insertion bonus $\beta$.
@@ VI Datasets
Our system is evaluated on three widely recognized ASR datasets, leveraging pre-trained models that require no additional training or data collection. The evaluation datasets comprise the Wall Street Journal (WSJ0) test subset, TED-LIUM Release 3 test set, and a set of native American English speakers from the ALLSSTAR corpus [40].
@@
From the ALLSSTAR corpus [40], we select recordings of 26 native English speakers, comprising a total of 3,060 utterances, to establish a controlled evaluation baseline aimed at consistent speech patterns.
@@ VII Experiments
Our experimental framework evaluates multiple acoustic and language models in a zero-shot setting, utilizing pre-trained models without any additional fine-tuning or adaptation. We utilized two state-of-the-art pre-trained acoustic models. The first is wav2vec 2.0 [9]. We evaluated two model sizes: Base (12 transformer encoder blocks with 95M parameters) and Large (24 blocks, 317M parameters). These models were pre-trained on a substantial amount of unsupervised data and subsequently fine-tuned on various supervised corpus sizes: 10 minutes, 100 hours, and 960 hours. The second acoustic model is HuBERT [10], which comes in two model sizes: Large (24 encoder blocks, 300M parameters) and X-Large (48 encoder blocks, 1B parameters). All acoustic models were used in their off-the-shelf configurations.
We investigated three LLMs as language models: GPT-2 [14], Falcon [41], and LLaMA 2 [15]. Note that each LLM has a different set of tokens, all of which are BPE-based [42]. Hence, to match the different token sets to the acoustic models, we filtered each set to include only tokens comprising English alphabetic characters, excluding numeric and special characters, to ensure compatibility with the acoustic model character set.
The pre-processing pipeline includes adding a 0.5-second silence period to each audio input, which extends the input context and empirically improves model generation stability. We then employed Silero Voice Activity Detector (VAD) version 5 [43] to trim irrelevant silence periods, optimizing the input for transcription. The parameters of Silero VAD included a start extension of 0.2 seconds and no end extension.
The decoding process utilizes beam search with beam size $b=5$ and considers the top $k=5,000$ candidates from the LLM’s predictions. We will discuss how these parameters were empirically determined in Section VII-D. The decoding process employs three distinct stopping criteria: (i) completion assessment, which evaluates the likelihood of a hypothesis to be a complete sentence against the probability of adding any another token; (ii) validation of acoustic probability threshold, which terminates paths where new token additions result in probabilities falling below a threshold of 0.3, excluding whitespace tokens; and (iii) audio length alignment, which stops decoding when the generated transcript aligns with the entire duration of the input audio. These criteria work together to ensure both transcription completeness and acoustic fidelity.
@@ VII-A Hyperparameters
TABLE II: ASR performance comparison across different acoustic and language models on ALLSSTAR Eng, WSJ0, and TED-LIUM 3 datasets. WER and CER values indicate lower is better ( $\downarrow$). Bold values are the best within an acoustic-model group; Values marked with ’’ are the overall best for a given metric. LLM beam rows correspond to our proposed method.
@@
VII-B Alignment Optimization
We implemented two main optimizations to improve computational efficiency. Firstly, we cached the end frame and alignment scores from prior steps in the beam search process to minimize redundant calculations. Instead of recalculating alignments from the beginning, each new step utilized the cached frame minus one position, combining the stored scores with new calculations to refine the results. Empirical testing demonstrated that achieving optimal alignment accuracy required a one-frame overlap. For the second optimization, we established a forward-looking boundary. Instead of aligning each prefix text with the entire audio sequence, we restricted the search to a maximum of 1,500 milliseconds (75 frames) ahead of the current position. These optimizations significantly reduced memory usage and computation time.
@@ VII-C Comparing different acoustic models and LLMs
Table II provides comprehensive performance comparisons across all model configurations and datasets. We evaluate four decoding baselines for each acoustic model. The first baseline, greedy, performs decoding without any language model. The second, beam 4-gram, uses beam search with a KenLM 4-gram language model, a beam size of 1500, language model weight 2.0, and word insertion penalty -1.0. The third, Transformer LM, utilizes a 20-layer Transformer-based language model with a beam size of 500, a language model weight of 2.0, and a word insertion penalty of –1.0. The fourth, GCNN, employs a convolutional language model with a beam size of 80, a beam threshold of 10, a language model weight of 1.0, and a word insertion bonus of 2.0. All decoding configurations follow the setups reported in [9, 44], using the Flashlight decoder via Fairseq’s interface [45, 46].
Our experiments demonstrate that combining the HuBERT X-Large acoustic model with the LLaMA 2 language model achieves the best performance across most datasets. This configuration yields the lowest WER and CER on WSJ0 (WER: 4.04%, CER: 0.72%) and TED-LIUM 3 (WER: 6.81%, CER: 2.73%). For the ALLSSTAR English dataset, the HuBERT Large with GPT-2 achieves optimal performance (WER: 4.57%, CER: 2.15%). These results align with the individual strengths of these models: HuBERT’s superior acoustic modeling capabilities in the speech domain and the advanced language understanding demonstrated by LLaMA 2 in their respective evaluations.
The impact of pre-training data volume on model performance is significant, as shown in II. Models trained on limited data (10-minute subset) consistently underperform across all decoding strategies, with WER increases of up to 31.35% for wav2vec 2.0 Base. However, as the pre-training data volume increases to 100 hours and 960 hours, our approach demonstrates substantial improvements over the baselines, highlighting the importance of sufficient pre-training data for effectively integrating large language models.
Interestingly, GCNN and Transformer LMs outperform our LLM-guided decoder on the ALLSSTAR dataset. This is likely due to the dataset’s characteristics—short, syntactically simple utterances with minimal vocabulary and limited context, which diminish the benefits of long-range reasoning. In contrast, on TED-LIUM 3, which includes longer utterances, named entities, and syntactic structure, our LLM-guided decoder consistently outperforms traditional LMs, highlighting its ability to model global context and semantic dependencies.
@@
[…]

### v2

@@ III-C Decoding
The decoding mechanism refers to the process of inferring the most probable sequence of words or tokens given an input speech signal. It combines the outputs of the acoustic model and the language model to generate the final transcription using an efficient search algorithm, such as beam search, stack decoding, or fast match [huang2001spoken]. Before introducing our decoding method, we briefly review decoding in HMM-based ASR and in SSL-ASR systems. This comparison provides context for evaluating our approach relative to the former, which is somewhat probabilistically principled (ignoring $\alpha$ and $\beta$), and the latter, which lacks a fully consistent probabilistic formulation.
@@
where $\alpha$ scales the influence of the language model, $\beta$ is the word insertion penalty that penalizes (or rewards) the number of words to avoid overgeneration or undergeneration, and |W||{W}| is the number of words in the hypothesis. This formulation allows a beam search or Viterbi decoder to efficiently explore the space of possible word sequences (as expressed by the HMM states) and choose the one with the highest combined score [rabiner1993fundamentals].
In contrast to HMMs, acoustic models that are based on wav2vec 2.0 [baevski2020wav2vec] and HuBERT [hsu2021hubert] are trained using the CTC loss function [graves2006connectionist] on characters. During decoding of these models, a prefix beam search decoding is used, and the scoring function for each candidate word sequence W{W} is formulated as
@@
where we assumed that the posterior probability over the characters $P(Y|X)$ can be easily translated into a probability over words $P_{\text{CTC}}({W}|\mathbf{X})$. Note that systems based on these acoustic models incorporate language models by multiplying the acoustic model $P_{\text{CTC}}({W}|\mathbf{X})$ by P(W)P({W}) [schneider2019wav2vec, baevski2020wav2vec]. This scoring function is the standard shallow-fusion formulation. While this approach yields high performance, it is implausible as it undermines the probabilistic meaning of the models.
@@ IV Proposed Approach
We propose integrating the LLM directly into the decoding process iteratively. At each step, the LLM proposes a set of candidate tokens, and each proposal is acoustically evaluated by matching it to the speech signal by using the SSL model and combining its alignment likelihood with the LLM probability inside the MAP recursion. This ensures that only candidates with high scores from both the LLM and acoustic models remain in the beam. Unlike post-processing approaches, which operate on a completed transcript, our method incorporates the LLM within the decoding process itself. As we will explain next, this integration is implemented efficiently using dynamic programming.
We propose integrating the LLM directly into the decoding process iteratively. At each step, the LLM proposes a set of candidate tokens, and each proposal is acoustically evaluated by matching it to the speech signal by using the SSL model and combining its alignment likelihood with the LLM probability inside the MAP recursion. This ensures that only candidates with high scores from both the LLM and acoustic models remain in the beam. Unlike post-processing approaches, which operate on a completed transcript, our method incorporates the LLM within the decoding process itself. Prior work [ma2024asr] identifies tokenizer and segmentation mismatches between ASR systems and LLMs as a challenge for integrated decoding; our alignment-based scoring provides a direct way to address this issue. As we will explain next, this integration is implemented efficiently using dynamic programming.
From this point onward, we will use tokens instead of words; thus, $w_{n}\in\mathcal{V}$ denotes a token from the set of all tokens $\mathcal{V}$. Define the alignment sequence $A=(a_{1},\ldots,a_{N})$, where $a_{n}\in[0,T-1]$ represents the start time of token $w_{n}$ for $1\leq n\leq N$. We denote by $\mathcal{A}({W})$ the set of all possible alignments (i.e., start times) corresponding to a given token sequence W{W}. Lastly, denote by $a_{1}^{n}=(a_{1},\ldots,a_{n})$ the alignment (start times) of the first $n$ tokens, $w_{1}^{n}=(w_{1},\ldots,w_{N})$.
Our objective is to derive a MAP estimate of the most likely token sequence, ${W}^{*}$. Rather than incorporating the probability P(W)P({W}) as a whole, we decompose it into a sequence of token-level predictions. At each step, we evaluate candidate tokens by combining their LLM likelihoods with an alignment-based acoustic score derived from the SSL emissions. More formally, we start from the MAP estimator given in (1):
@@
We approximate this expression by assuming that the most likely alignment dominates the total probability. That is, instead of summing over all valid alignments, we approximate this sum by the alignment with the highest posterior probability,
${W}^{*}\simeq\arg\max_{{W}}\max_{A\in\mathcal{A}({W})}P({W},A\mid\mathbf{X})~,$ (9)
following the same approximation used in the original CTC decoding formulation [pmlr-v32-graves14].
This probability can be decomposed into two terms: a term that is associated with the acoustic model and a term that is related to the language model. We start by expressing the sequences W{W} and $A$ explicitly:
@@
In summary, we reformulated the MAP estimator as the sum of two terms corresponding to the acoustic and language models. The acoustic model term is expressed as $\sum_{n}P(a_{n}|w_{n},a_{1}^{n-1},w_{1}^{n-1},\mathbf{X})$, in contrast to the conventional forms $P(\mathbf{X}|{W})$ or $P_{\text{CTC}}({W}|\mathbf{X})$. The language model term is given by $\sum_{n}P(w_{n}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$, as opposed to the standard formulation $P({W})=\sum_{n}P(w_{n}|w_{1}^{n-1})$. It is important to note that these terms inherently assume access to the entire speech signal $\mathbf{X}$, which limits the applicability of the proposed method in streaming scenarios. A simple workaround is to process speech in asynchronous chunks, allowing for partial processing. While more sophisticated strategies could be developed to fully enable streaming functionality, such approaches are beyond the scope of this work.
@@ V Algorithmic Implementation
We proceed to show that the MAP estimate in (10) can be evaluated iteratively, allowing the probabilities to be computed incrementally at each step. Equation (11) defines the incremental score of a single partial hypothesis. In practice, the decoder applies this update to all hypotheses maintained in a beam of size $B$, and keeps only the highest-scoring extensions at each step (see Algorithm 1).
Suppose we have already computed the probability up to token n1n-1, that is, we know $\log P(w_{1}^{n-1},a_{1}^{n-1}|\mathbf{X})$. The probability of token $n$ is obtained by incrementally adding two terms: the language model probability, $\log P(w_{n}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$ and the acoustic alignment probability $\log P(a_{n}|w_{n},a_{1}^{n-1},w_{1}^{n-1},\mathbf{X})$. Overall, we have,
@@
The iterative procedure operates as follows. At step $n$, we prompt the LLM with the sequence $w_{1}^{n-1}$ (already aligned to the speech) and sample $K$ candidate tokens $\{w_{n}^{(k)}\}_{k=1}^{K}$, each with an associated probability $P(w_{n}^{(k)}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$. Each candidate token is then aligned to the speech signal $\mathbf{X}$, assuming the last token began at $a_{n-1}$. The alignment is performed using the Viterbi algorithm [kurzinger2020ctc], which computes the most likely alignment of each token’s characters to the CTC emissions [^1] For example, this can be implemented by performing forced alignment of the tokens $w_{n}$ to the CTC emission outputs of wav2vec 2.0 or HuBERT.. For each candidate, we compute the combined probability of the token and its alignment, namely $P(w_{n}|w_{1}^{n-1},a_{1}^{n-1},\mathbf{X})$ and $P(a_{n}|w_{n},a_{1}^{n-1},w_{1}^{n-1},\mathbf{X})$, which together define the incremental score in Equation (11) and Equation (12). In implementation, this update is applied to all partial hypotheses in the beam, and the top $B$ extensions are retained (see Algorithm 1).
@@
Algorithm 1 Zero-Shot LLM-Driven ASR Decoder
@@
18:Return: $\textsc{BEAM.TopB}(\mathcal{B}_{n},1)$ {best beam defines $\mathbf{W}^{*}$}
We conclude this section by summarizing the proposed method in the form of pseudo-code, presented in Algorithm 1. The algorithm takes as input a speech signal $\mathbf{X}$, an acoustic model (wav2vec 2.0 [baevski2020wav2vec] or HuBERT [hsu2021hubert]) denoted as AM, and a large language model, (e.g., LLaMA 2, Falcon or GPT-2), denoted as LLM. Additional input parameters include the beam size $B$, the number of candidate tokens sampled at each iteration from the LLM $K$, the LLM scaling factor $\alpha$, and the token insertion bonus $\beta$.
@@ VI Datasets
Our system is evaluated on three widely recognized ASR datasets, leveraging pre-trained models that require no additional training or data collection. The evaluation datasets comprise the Wall Street Journal (WSJ0) test subset, TED-LIUM Release 3 test set, and a set of native American English speakers from the ALLSSTAR corpus [Bradlow_ALLSSTAR].
@@
From the ALLSSTAR corpus [Bradlow_ALLSSTAR], we select recordings of 26 native English speakers, comprising a total of 3,060 utterances, to establish a controlled evaluation baseline aimed at consistent speech patterns.
@@ VII Experiments
Our experimental framework evaluates multiple acoustic and language models in a zero-shot setting, utilizing pre-trained models without any additional fine-tuning or adaptation. We utilized two state-of-the-art pre-trained acoustic models. The first is wav2vec 2.0 [baevski2020wav2vec]. We evaluated two model sizes: Base (12 transformer encoder blocks with 95M parameters) and Large (24 blocks, 317M parameters). These models were pre-trained on a substantial amount of unsupervised data and subsequently fine-tuned on various supervised corpus sizes: 10 minutes, 100 hours, and 960 hours. The second acoustic model is HuBERT [hsu2021hubert], which comes in two model sizes: Large (24 encoder blocks, 300M parameters) and X-Large (48 encoder blocks, 1B parameters). All acoustic models were used in their off-the-shelf configurations.
We investigated three LLMs as language models: GPT-2 [radford2019language], Falcon [penedo2023refinedweb], and LLaMA 2 [touvron2023llama]. Note that each LLM has a different set of tokens, all of which are BPE-based [sennrich-etal-2016-neural]. Hence, to match the different token sets to the acoustic models, we filtered each set to include only tokens comprising English alphabetic characters, excluding numeric and special characters, to ensure compatibility with the acoustic model character set.
The pre-processing pipeline includes adding a 0.5-second silence period to each audio input, which extends the input context and empirically improves model generation stability. We then employed Silero Voice Activity Detector (VAD) version 5 [SileroVAD] to trim irrelevant silence periods, optimizing the input for transcription. The parameters of Silero VAD included a start extension of 0.2 seconds and no end extension.
The decoding process utilizes beam search with beam size $b=5$ and considers the top $k=5,000$ candidates from the LLM’s predictions. We will discuss how these parameters were empirically determined in Section VII-D. To ensure compatibility between the acoustic emissions and the language model, we restrict candidate tokens to those that fall within the acoustic model’s character set. The decoder process employs three distinct stopping criteria: (i) completion assessment, which evaluates the likelihood of a hypothesis to be a complete sentence against the probability of adding any another token; (ii) validation of acoustic probability threshold, which terminates paths where new token additions result in probabilities falling below a threshold of 0.3, excluding whitespace tokens; and (iii) audio length alignment, which stops decoding when the generated transcript aligns with the entire duration of the input audio. These criteria work together to ensure both transcription completeness and acoustic fidelity.
@@ VII-A Hyperparameters
TABLE II: ASR performance comparison across different acoustic and language models on ALLSSTAR Eng, WSJ0, and TED-LIUM 3 datasets. WER and CER values indicate lower is better ( $\downarrow$). Bold values are the best within an acoustic-model group; values marked with ’’ are the overall best for a given metric. LLM beam rows correspond to our proposed method. All LM-based beam-search baselines use the standard shallow-fusion formulation.
@@
VII-B Algorithmic Optimizations
We implemented several optimizations to improve computational efficiency. To accelerate iterative decoding, we reuse the LLM’s cached key-value $(K,V)$ states as the prompt evolves, avoiding repeated full-prompt attention computations and reducing both runtime and memory usage.
We also implemented two main optimizations for the acoustic alignment process. Firstly, we cached the end frame and alignment scores from prior steps in the beam search process to minimize redundant calculations. Instead of recalculating alignments from the beginning, each new step utilized the cached frame minus one position, combining the stored scores with new calculations to refine the results. Empirical testing demonstrated that achieving optimal alignment accuracy required a one-frame overlap. For the second optimization, we established a forward-looking boundary. Instead of aligning each prefix text with the entire audio sequence, we restricted the search to a maximum of 1,500 milliseconds (75 frames) ahead of the current position. These optimizations significantly reduced memory usage and computation time.
@@ VII-C Comparing different acoustic models and LLMs
Table II provides comprehensive performance comparisons across all model configurations and datasets. We evaluate four decoding baselines for each acoustic model. The first baseline, greedy, performs decoding without any language model. The second, beam 4-gram, uses beam search with a KenLM 4-gram language model, a beam size of 1500, language model weight 2.0, and word insertion penalty -1.0. The third, Transformer LM, utilizes a 20-layer Transformer-based language model with a beam size of 500, a language model weight of 2.0, and a word insertion penalty of -1.0. The fourth, GCNN, employs a convolutional language model with a beam size of 80, a beam threshold of 10, a language model weight of 1.0, and a word insertion bonus of 2.0. All LM-based beam-search baselines use the standard shallow-fusion formulation, following the configurations in [baevski2020wav2vec, hannun2019tds] and using the Flashlight decoder via Fairseq’s interface [pratap2019wav2letter, ott2019fairseq].
Our experiments demonstrate that combining the HuBERT X-Large acoustic model with the LLaMA 2 language model achieves the best performance across most datasets. This configuration yields the lowest WER and CER on WSJ0 (WER: 4.04, CER: 0.72) and TED-LIUM 3 (WER: 6.81, CER: 2.73). For the ALLSSTAR English dataset, the HuBERT Large with GPT-2 achieves optimal performance (WER: 4.57, CER: 2.15). These results align with the individual strengths of these models: HuBERT’s superior acoustic modeling capabilities in the speech domain and the advanced language understanding demonstrated by LLaMA 2 in their respective evaluations.
The impact of pre-training data volume on model performance is significant, as shown in II. Models trained on limited data (10-minute subset) consistently underperform across all decoding strategies, with WER increases of up to 31.35 for wav2vec 2.0 Base. However, as the pre-training data volume increases to 100 hours and 960 hours, our approach demonstrates substantial improvements over the baselines, highlighting the importance of sufficient pre-training data for effectively integrating large language models.
Interestingly, GCNN and Transformer LMs outperform our LLM-guided decoder on the ALLSSTAR dataset. This is likely due to the dataset’s characteristics-short, syntactically simple utterances with minimal vocabulary and limited context, which diminish the benefits of long-range reasoning. In contrast, on TED-LIUM 3, which includes longer utterances, named entities, and syntactic structure, our LLM-guided decoder consistently outperforms traditional LMs, highlighting its ability to model global context and semantic dependencies.
@@
Figure 2: WER and CER using HuBERT XLarge with LLaMA 2 for different top-$k$ values, where $k$ represents the number of most probable next-token candidates the LLM considers for each prefix during transcription. Results are shown for WSJ0, ALLSSTAR English, and TED-LIUM datasets.